OpenStudy (anonymous):

Together, Bob and Tom take 4 hours to get the yard work done. If Bob works alone, it takes him 6 hours. How long would it take Tom, working alone, to do all of the yard work? 10 hours 12 hours 8 hours 9 hours

hero (hero):

B = hours Bob can complete job alone M = hours Tom can complete job alone T = Time it takes if they both work together Working Together Formula: \[\frac{BM}{B + M} = T\] B = 6 T = 4 M = ? \[\frac{6M}{6 + M} = 4\]

hero (hero):

Solve for \(M\)

hero (hero):

So what does M equal?